# Personalized Enhanced Federated Multi-View Clustering via Heat-Kernel Tensor Decomposition

### v2

@@ Abstract
This paper proposes a personalized federated learning framework integrating heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with tensor decomposition techniques. The approach combines heat-kernel coefficients adapted from quantum field theory with PARAFAC2 and Tucker decomposition to transform distance metrics and efficiently represent high-dimensional multi-view structures. Two main algorithms, FedHK-PARAFAC2 and FedHK-Tucker, are developed to extract shared and view-specific features while preserving inter-view relationships. The framework addresses data heterogeneity, privacy preservation, and communication efficiency challenges in federated learning environments. Theoretical analysis provides convergence guarantees, privacy bounds, and complexity analysis. The integration of heat-kernel methods with tensor decomposition in a federated setting offers a novel approach for effective multi-view data analysis while ensuring data privacy.
Keywords: Federated learning, Multi-view clustering, Heat-kernel methods, Tensor decomposition, PARAFAC2, Tucker decomposition, Privacy preservation.
1 Introduction
@@
Tensor-based machine learning approaches have emerged as powerful tools for analyzing multi-dimensional data structures, enabling the capture of complex relationships and interactions across multiple modes. Tensors, as generalizations of matrices to higher dimensions, provide a natural framework for representing multi-view data, where each view corresponds to a different modality or feature set. For instance, in image processing, a color image can be represented as a third-order tensor with dimensions corresponding to height, width, and color channels (Red, Green, Blue) as illustrated in Figure 1.
Figure 1: The Illustration of A 3-Dimensional of a Red-Green-Blue (RGB) Image
As can be seen in Figure 1, a color image is represented as a 3-dimensional tensor, where each slice along the third dimension corresponds to one of the RGB color channels. This tensor representation allows for the preservation of spatial and color information, enabling more effective analysis and processing of image data. RGB image is stored in a 3-Dimensional array as $\text{channels}\times\text{height}\times\text{width}$ or $\text{height}\times\text{width}\times\text{channels}$. The $7\times 7\times 3$ tensor in Figure 1 represents a small RGB image with a height and width of 7 pixels and 3 color channels. Each element in the tensor corresponds to the intensity value of a specific color channel at a particular pixel location.
@@
This paper presents two novel algorithms, namely FedHK-PARAFAC2 and FedHK-Tucker, that integrate heat-kernel enhanced tensorized multi-view fuzzy c-means clustering with advanced tensor decomposition techniques in a federated learning framework. The proposed methods leverage heat-kernel coefficients adapted from quantum field theory to transform conventional distance metrics, enabling the capture of local geometric relationships within multi-view data distributed across heterogeneous clients. By incorporating PARAFAC2 and Tucker decomposition, the algorithms efficiently represent high-dimensional multi-view structures while addressing challenges of data heterogeneity, privacy preservation, and communication efficiency inherent in federated learning environments.
@@ 2.1 Federated Learning and Clustering
Pattern recognition and clustering in federated learning (FL) environments have garnered significant attention due to the increasing need for privacy-preserving distributed data analysis. Federated learning, introduced by McMahan et al. [mcmahan2017communication], enables collaborative model training across decentralized clients without sharing raw data. While initial research focused on supervised learning tasks, recent efforts have extended federated learning to unsupervised clustering scenarios. Early work by Dennis et al. [dennis2021heterogeneity] explored federated k-means clustering, demonstrating the feasibility of distributed clustering with local updates and global aggregation. Ghosh et al. [ghosh2002consensus] proposed consensus-based approaches for federated hierarchical clustering, addressing challenges of data heterogeneity and communication efficiency. Makhija et al. [makhija2022federated] introduced federated fuzzy c-means clustering, leveraging soft assignments to enhance robustness against noise and outliers. However, these methods often assume homogeneous data distributions and do not fully exploit multi-view data structures prevalent in modern applications. Recent advancements have focused on personalized federated clustering to accommodate client heterogeneity. Caldarola et al. [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. Privacy-preserving aspects have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical], ensuring that sensitive information remains protected during the clustering process. Communication efficiency has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq], reducing the overhead associated with transmitting model updates across clients.
@@ 2.2 Multi-View Clustering
Multi-view clustering (MVC) has emerged as a powerful paradigm for leveraging complementary information from multiple data representations to enhance clustering performance. Early work by Bickel and Scheffer [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating that integrating multiple views can lead to improved clustering accuracy. Classical MVC approaches can be categorized into co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as those proposed by Kumar et al. [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering. Graph-based MVC has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. Deep learning approaches have revolutionized MVC through end-to-end representation learning. Zhang et al. [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. [jiang2014collaborative] and Wang et al. [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. Sinaga et al. [sinaga2025parameter] proposed a parameter-free multi-view fuzzy clustering algorithm that automatically learns view weights and fuzzification parameters, enhancing clustering robustness and interpretability. However, these methods typically require centralized data access and are not designed for federated environments.
@@
The federated multi-view clustering is first introduced by Chen et al. [chen2022federated] for federated multi-view representation learning and Liu et al. [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication. Yang et al. [yang2024federated] proposed a federated multi-view clustering framework that combines local clustering with global consensus, but it lacks the integration of heat-kernel methods and advanced tensor decomposition techniques. Earlier, Sinaga [sinaga2025fedhk] proposed a heat-kernel enhanced federated multi-view clustering framework, but it did not incorporate tensor decomposition methods for efficient multi-view representation. Thus, there is a need for advanced federated multi-view clustering approaches that leverage tensor decomposition and heat-kernel methods to address data heterogeneity, privacy preservation, and communication efficiency challenges in distributed settings.
@@ 2.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds. The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis. In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum]. The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
@@ 2.4 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary N-way data structures. Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems. Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries. Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors. The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power. In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Zhou et al. [cichocki2015tensor] developed tensor-based multiway arrays for bid data sets, while Benjamin et al. [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.

### v3

@@ Abstract
This paper introduces mathematical frameworks that address the challenges of multi-view clustering in federated learning environments. The objective is to integrate optimization techniques based on new objective functions employing heat-kernel coefficients to replace conventional distance metrics with quantum-inspired measures. The proposed frameworks utilize advanced tensor decomposition methods—specifically, PARAFAC2 and Tucker decomposition—to efficiently represent high-dimensional, multi-view data while preserving inter-view relationships. The research has yielded the development of four novel algorithms: an efficient federated kernel multi-view clustering (E-FKMVC) model; FedHK-PARAFAC2; FedHK-Tucker; and FedHK-MVC-Person with PARAFAC2 Decomposition (Personalized FedHK-PARAFAC2). The primary objective of these algorithms is to enhance the efficacy of clustering processes while ensuring the confidentiality and efficient communication in federated learning environments. Theoretical analyses of convergence guarantees, privacy bounds, and complexity are provided to validate the effectiveness of the proposed methods. In essence, this paper makes a significant academic contribution to the field of federated multi-view clustering through its innovative integration of mathematical modeling and algorithm design. This approach addresses the critical challenges of data heterogeneity and privacy concerns, paving the way for enhanced data management and analytics in various contexts.
Keywords: Federated learning, Multi-view clustering, Heat-kernel methods, Tensor decomposition, PARAFAC2, Tucker decomposition, Privacy preservation, E-FKMVC, FedHK-PARAFAC2, FedHK-Tucker, Personalized FedHK-PARAFAC2/Tucker
@@
Tensor-based machine learning approaches have emerged as powerful tools for analyzing multi-dimensional data structures, enabling the capture of complex relationships and interactions across multiple modes. Tensors, as generalizations of matrices to higher dimensions, provide a natural framework for representing multi-view data, where each view corresponds to a different modality or feature set. For instance, in large language models (LLMs), text data can be represented as a three-dimensional tensor, with dimensions corresponding to words, contexts, and semantic features. This tensor representation allows for the preservation of contextual information and facilitates the extraction of meaningful patterns through tensor decomposition techniques. In computer vision, images and videos can be represented as higher-order tensors, capturing spatial and temporal relationships across multiple channels. For example, a color image can be represented as a third-order tensor with dimensions corresponding to height, width, and color channels (RGB). This tensor representation enables the preservation of spatial and color information, facilitating more effective analysis and processing of image data. As illustrated in Figure 1, a color image is represented as a third-order tensor, where each slice along the third dimension corresponds to one of the RGB color channels. This tensor representation allows for the preservation of spatial and color information, enabling more effective analysis and processing of image data.
Figure 1: The Illustration of a Color Image as a Third-Order Tensor Representation with RGB Channels. The channel dimension captures the color information, while the height and width dimensions represent the spatial structure of the image.
As demonstrated in Figure 2, multi-view data can be effectively represented as higher-order tensors, capturing the complex relationships and interactions across different views. Each view corresponds to a different modality or feature set, and the tensor representation allows for the preservation of inter-view relationships. A color image dataset with multiple views, such as RGB, depth, infrared (IR), and thermal channels, can be represented as a fourth-order tensor. Each slice along the fourth dimension corresponds to a different view of the same image, capturing complementary information across modalities. This tensor representation enables the application of advanced tensor decomposition techniques to extract shared and view-specific features, facilitating more effective multi-view data analysis.
WidthHeightViewsView 1 (RGB)View 2 (Depth)View 3 (IR)View 4 (Thermal)4th mode
Figure 2: Tensor Representation of Multi-View Data with Different Modalities. Each slice along the fourth dimension corresponds to a different view of the same image, capturing complementary information across modalities such as RGB, depth, infrared (IR), and thermal channels.
@@
This paper presents new frameworks that integrate heat-kernel coefficients and multi-view fuzzy c-means clustering via advanced tensor decomposition techniques. The framework’s scope extends beyond the realm of addressing challenges related to data heterogeneity, privacy preservation, and communication efficiency in federated learning environments. Furthermore, it overcomes the limitations of existing MV clustering methods in both federated and non-federated settings, thereby addressing the issue of data decomposition in MV clustering that does not employ deep learning. The contribution of this paper is threefold.
  • Four novel algorithms are presented, based on a basic, efficient federated kernel multi-view clustering (E-FKMVC) model. This model integrates the heat-kernel method to replace the conventional Euclidean distance metric with heat-kernel coefficients adapted from quantum field theory. The FedHK-PARAFAC2, FedHK-Tucker, and Personalized FedHK-PARAFAC2/Tucker algorithms are based on the E-FKMVC objective function.
  • The integration of tensor decomposition techniques, specifically PARAFAC2 and Tucker decomposition, into the federated multi-view clustering framework to efficiently represent high-dimensional multi-view structures while preserving inter-view relationships.
  • Theoretical analysis of the proposed algorithms, including convergence guarantees, privacy bounds, and complexity analysis, to ensure their effectiveness and efficiency in federated learning environments.
The subsequent sections of this paper are organized as follows. In section 1, the extant literature on federated learning and clustering, multi-view clustering, heat kernel methods and quantum-inspired learning, and tensor decomposition techniques is reviewed. In Section 2, the reader will find concise theoretical and mathematical backgrounds on federated learning and clustering, multi-view clustering, heat kernel methods, and tensor decomposition. The proposed efficient federated kernel multi-view clustering (E-FKMVC) framework is delineated in Section 3. As delineated in Section 4, the proposed FedHK-PARAFAC2, FedHK-Tucker, and its personalized algorithms are described. In accordance with the findings of the research, Section 5 of the paper presents a conclusion. Additionally, it delineates the potential avenues for future investigation and the practical applications of the research.
@@ 1.1 Federated Learning and Clustering
Pattern recognition and clustering in federated learning (FL) environments have garnered significant attention due to the increasing need for privacy-preserving distributed data analysis. Generally, federated learning can be categorized into two types: horizontal federated learning (HFL) and vertical federated learning (VFL) (See Riviera et al. (2025) [11000331] for a comprehensive survey on FL). HFL is applicable when datasets across clients share the same feature space but differ in samples, while VFL is suitable when datasets share the same sample space but differ in features. Early work by McMahan et al. (2017) [mcmahan2017communication] introduced the Federated Averaging (FedAvg) algorithm, which laid the groundwork for subsequent FL research. Konečnỳ et al. (2016) [konevcny2016federated] introduced federated optimization as a method to train a centralized model using decentralized data stored across numerous devices, emphasizing communication efficiency due to limited connectivity and privacy concerns. They proposed modifications to existing algorithms to adapt them to federated optimization settings. Geyer et al. (2017) [geyer2017differentially] demonstrated how differential privacy can be integrated into federated clustering while maintaining acceptable utility. Li et al. (2020) [li2020acceleration] adopted the foundational ideas of federated optimization and extended them by integrating advanced techniques like gradient compression, acceleration, and variance reduction to address communication bottlenecks and improve the scalability of federated learning systems. However, these privacy mechanisms often come at the cost of accuracy, particularly in high-dimensional settings. In the context of clustering, several approaches have been proposed to adapt traditional clustering algorithms to federated settings. For instance, Dennis et al. (2021) [dennis2021heterogeneity] explored federated k-means clustering, demonstrating the feasibility of distributed clustering with local updates and global aggregation. Ghosh et al. (2002) [ghosh2002consensus] proposed consensus-based approaches for federated hierarchical clustering, addressing challenges of data heterogeneity and communication efficiency. Makhija et al. (2022) [makhija2022federated] introduced federated fuzzy c-means clustering, leveraging soft assignments to enhance robustness against noise and outliers. However, these methods often assume homogeneous data distributions and do not fully exploit multi-view data structures prevalent in modern applications. Recent advancements have focused on personalized federated clustering to accommodate client heterogeneity. Caldarola et al. (2021) [caldarola2021cluster] proposed adaptive weight schemes for heterogeneous clustering, while Li et al. (2022) [li2022hpfl] developed meta-learning approaches for personalized cluster initialization. Privacy-preserving aspects have been addressed through differential privacy mechanisms [kairouz2021advances] and secure aggregation protocols [bonawitz2017practical], ensuring that sensitive information remains protected during the clustering process. Communication efficiency has been tackled through gradient compression [konecny2016federated], sparsification techniques [aji2017sparse], and model quantization [reisizadeh2020fedpaq], reducing the overhead associated with transmitting model updates across clients. However, these methods often face challenges related to data heterogeneity, communication efficiency, and convergence guarantees in distributed environments.
@@ 1.2 Multi-View Clustering
Multi-view clustering (MVC) has emerged as a powerful paradigm for leveraging complementary information from multiple data representations to enhance clustering performance. Early work by Bickel and Scheffer (2004) [bickel2004multi] established the theoretical foundations for multi-view learning, demonstrating that integrating multiple views can lead to improved clustering accuracy. Classical MVC approaches can be categorized into co-training based methods, subspace learning approaches, and consensus clustering techniques. Co-training methods, such as those proposed by Kumar et al. (2011) [kumar2011co], iteratively refine clustering results by leveraging inter-view consistency. Subspace learning approaches, including Canonical Correlation Analysis (CCA) based methods [hotelling1992relations] and shared subspace clustering [wang2019multiple], project multiple views into a common representation space for unified clustering. Graph-based MVC has gained significant attention due to its ability to capture complex inter-view relationships. Wang et al. (2019) [wang2019gmc] proposed multi-view graph clustering with adaptive neighbor learning, while Nie et al. (2016) [nie2016parameter] developed parameter-free auto-weighted multiple graph learning. Deep learning approaches have revolutionized MVC through end-to-end representation learning. Zhang et al. (2019) [zhang2019ae2] introduced auto-encoder based multi-view clustering, while Zhu et al. (2019) [zhu2019multi] developed deep multi-view clustering networks with view correlation discovery. Fuzzy clustering extensions to multi-view scenarios have been explored by Jiang et al. (2014) [jiang2014collaborative] and Wang et al. (2017) [wang2017multi], introducing weighted aggregation schemes and view-specific fuzzification parameters. Sinaga et al. (2025) [sinaga2025parameter] proposed a parameter-free multi-view fuzzy clustering algorithm that automatically learns view weights and fuzzification parameters, enhancing clustering robustness and interpretability. However, these methods typically require centralized data access and are not designed for federated environments.
@@
The federated multi-view clustering is first introduced by Chen et al. (2022) [chen2022federated] for federated multi-view representation learning and Liu et al. (2023) [liu2023federated] for privacy-preserving multi-view clustering. However, these methods do not leverage tensor decomposition for efficient representation and communication. Yang et al. (2024) [yang2024federated] proposed a federated multi-view clustering framework that combines local clustering with global consensus, but it lacks the integration of heat-kernel methods and advanced tensor decomposition techniques. Earlier, Sinaga (2025) [sinaga2025fedhk] proposed a heat-kernel enhanced federated multi-view clustering framework, but it did not incorporate tensor decomposition methods for efficient multi-view representation. Thus, there is a need for advanced federated multi-view clustering approaches that leverage tensor decomposition and heat-kernel methods to address data heterogeneity, privacy preservation, and communication efficiency challenges in distributed settings.
@@ 1.3 Heat Kernel Methods and Quantum-Inspired Learning
Heat kernel methods have emerged as powerful tools for capturing local geometric structure in high-dimensional data spaces. Originally developed in differential geometry and quantum field theory, heat kernels provide a natural way to encode multi-scale geometric information through diffusion processes on manifolds. The heat kernel, defined as the fundamental solution to the heat equation, has found extensive applications in machine learning through kernel methods and manifold learning. Ham et al. (2004) [ham2004kernel] introduced diffusion maps that leverage heat kernel properties for dimensionality reduction, while Coifman and Lafon (2006) [coifman2006diffusion] established theoretical foundations for diffusion-based data analysis. In clustering applications, heat kernel methods have been used to capture local neighborhood structure and improve cluster boundary detection. Zelnik-Manor and Perona (2004) [zelnik2004self] developed self-tuning spectral clustering using local heat kernel scales, while von Luxburg (2007) [von2007tutorial] provided comprehensive analysis of heat kernel properties in graph-based clustering.
Quantum-inspired machine learning has gained attention for its ability to leverage quantum mechanical principles in classical algorithms. Quantum kernels, introduced by Havlíček et al. (2019) [havlivcek2019supervised], enable the computation of quantum feature maps for enhanced pattern recognition. The connection between quantum field theory and machine learning has been explored through path integral formulations [carleo2017solving] and quantum neural networks [biamonte2017quantum]. The integration of heat kernel coefficients with clustering algorithms represents a natural extension of kernel methods to soft clustering scenarios. By incorporating heat diffusion properties, clustering algorithms can better capture local data manifold structure and improve robustness to noise and outliers.
@@ 1.4 Tensor Decomposition Methods
Tensor decomposition has emerged as a powerful paradigm for analyzing high-dimensional multi-modal data, providing compact representations that capture multilinear relationships inherent in complex datasets. The field has evolved from classical matrix factorization techniques to sophisticated tensor algebra frameworks capable of handling arbitrary NwayN-way data structures. Canonical polyadic decomposition (CPD) or parallel factors (PARAFAC) [harshman1994parafac, carroll1970analysis], Tucker decomposition (TD), and non-negative tensor factorization (NTF) are well-known approaches to reorder a tensor into a matrix or vector. PARAFAC or CPD holds multilinearity assumptions such as handle a sum of rank-one tensors by targeting the same lengths. In this sense, the original PARAFAC decomposition can reveal uncorrelated (orthogonal) and correlated (oblique) factors in system-variation data matrices, but failed to describe correlated factors when applied to the cross-product matrices computed from one data [harshman1972parafac2, kiers1999parafac2]. PARAFAC2, on the other hand has the advantageous to handle irregular tensors by relaxing the canonical polyadic (CP) model [harshman1972parafac2]. PARAFAC2 enables the three-mode model, often implemented into subspace clustering algorithms taking advantageous of correlations between pair of factors. The PARAFAC2 normalize the generalization of PARAFAC to deal both with uncorrelated and correlated factors. Canonical Polyadic Decomposition (CPD), also known as CANDECOMP/PARAFAC, represents the foundational tensor decomposition method. Introduced independently by Carroll and Chang (1970) [carroll1970analysis] and Harshman [harshman1970foundations], CPD decomposes a tensor into a sum of rank-one tensors, providing a unique factorization under mild conditions. Tucker decomposition, proposed by Tucker [tucker1966some], provides a more flexible alternative through its core tensor and factor matrices structure. Unlike CPD, Tucker decomposition allows different ranks along each mode, making it suitable for data with varying complexity across dimensions. The Higher-Order SVD (HOSVD) algorithm by De Lathauwer et al. [de2000multilinear] established efficient computational procedures for Tucker decomposition, enabling its application to large-scale problems.
Recent advances in tensor decomposition have focused on addressing practical challenges such as missing data, noise robustness, and computational scalability. Tensor completion methods, including work by Liu et al. [liu2012tensor] and Yuan and Zhang [yuan2016tensor], enable handling of incomplete tensors through low-rank assumptions. Robust tensor decomposition techniques [huang2014robust] provide resilience against outliers and corrupted entries. Non-negative tensor factorization (NTF) has gained attention for applications requiring interpretable, parts-based representations. Lee and Seung’s pioneering work [lee2000algorithms] on non-negative matrix factorization was extended to tensors by Shashua and Hazan [shashua2005non], enabling decompositions with non-negativity constraints that often yield more interpretable factors. The integration of tensor decomposition with machine learning has led to tensorized neural networks [novikov2015tensorizing], tensor-based kernel methods [signoretto2011kernel], and multi-task learning frameworks [wimalawarne2014multitask]. These approaches leverage tensor structure to reduce model complexity while maintaining representational power. In the context of clustering, tensor-based methods have shown particular promise for multi-view and multi-way data analysis. Cichocki et al. (2015) [cichocki2015tensor] developed tensor-based multiway arrays for big data sets, while Benjamin et al. (2024) [benjamin2024tensor] proposed tensorized possibilistic c-means for high-dimensional clustering. However, these methods are not designed for federated environments and lack the privacy-preserving mechanisms required for distributed deployment.